### v2

# On an integrable discretization of the massive Thirring model in light-cone coordinates and the associated Yang–Baxter map

Abstract
We propose a fully discrete analog of the massive Thirring model in light-cone coordinates by constructing its Lax-pair representation. This Lax-pair representation can also be used to define a new Yang–Baxter map, so we obtain a Yang–Baxter map that admits a continuous limit. We present most of the results for the general case where the dependent variables are matrix-valued.
1 Introduction
In the early 1990s, V. G. Drinfeld [1] suggested the problem of studying the set-theoretical solutions to the quantum Yang–Baxter equation:
$R_{12}R_{13}R_{23}=R_{23}R_{13}R_{12},$ (1.1)
where $R_{ij}$ acts as a map $X\times X\to X\times X$ on the $i$th and $j$th factors of the Cartesian product $X\times X\times X$ and identically on the remaining factor, and $X$ is a set.
Veselov [2, 3] suggested to use the shorter term ‘Yang–Baxter map’ instead of ‘set-theoretical solution to the quantum Yang–Baxter equation’, which is now more common, so we also adopt it here. Since the suggestion by Drinfeld [1], the study of Yang–Baxter maps has attracted much attention and now the number of known Yang–Baxter maps is enormous and the results on Yang–Baxter maps appear to be somewhat scattered, except for some systematic classification results (see, e.g., [4]). This implies that the original problem setting by Drinfeld [1] would be too general, depending on the nature of the set $X$, and there exist too many (sometimes infinitely many) solutions to the quantum Yang–Baxter equation (1.1). Thus, it would be more realistic and meaningful to consider narrower versions of the Drinfeld problem, e.g., by restricting the admissible form of the map $R$ in a natural manner.
In this paper, we aim to construct a Yang–Baxter map that admits a natural continuous limit. In the simplest case, such a map $R$ involves an arbitrary (typically small) parameter $\hbar$ and takes the following form [5]:
$R=I+\hbar\hskip 1.0ptr+\mathcal{O}(\hbar^{2}),$ (1.2)
where $I$ is the identity mapping. The asymptotic expansion (1.2) with respect to $\hbar$ is a natural ansatz in that the quantum Yang–Baxter equation (1.1) for a linear operator $R$ reduces in the quasi-classical limit $\hbar\to 0$ to the classical Yang–Baxter equation [6, 7]:
$[r_{12},r_{13}]+[r_{12},r_{23}]+[r_{13},r_{23}]=0,$
where [,][\;\,,\;\,][ , ] is the commutator.
$q$$\widehat{q}$$\widehat{u}$$u$$R$
More specifically, if $q,u\in X$ and
$R:(q,u)\to\left(\widehat{q},\widehat{u}\right):=(q,u)+\hbar\hskip 1.0pt\left(f(q,u),g(q,u)\right)+\mathcal{O}(\hbar^{2}),$
the map $R$ provides in the continuous limit $\hbar\to 0$ a continuous system in $1+1$ space-time dimensions:
$\left\{\begin{split}&q_{t}=f(q,u),\\ &u_{x}=-g(q,u),\end{split}\right.$
where the subscripts $t$ and $x$ denote the partial differentiation with respect to these variables. In our previous paper [8], we already constructed a Yang–Baxter map of this type (see also some relevant results in [9, 10, 11, 12, 13, 14]).
We can consider a more general situation where the map $R_{12}$ admits the expansion in terms of (typically small) parameters $h$ and $\varDelta$ as
$R_{12}:(q,u)\to\left(\widehat{q},\widehat{u}\right):=(q,u)+\left(hf(q,u),\varDelta g(q,u)\right)+(\mathcal{O}(h^{2},h\varDelta),\mathcal{O}(h\varDelta,\varDelta^{2})).$ (1.3)
In this paper, we propose a Yang–Baxter map that admits the expansion (1.3) and provides in the continuous limit $h\to 0$ or/and $\varDelta\to 0$ a nontrivial semi-discrete/continuous system in $1+1$ space-time dimensions. The Yang–Baxter map proposed in this paper appears to be more general and interesting than the Yang–Baxter map constructed in our previous paper [8]; closely related results were obtained previously by Xu and Pelinovsky [15], but they did not explicitly present their results in the form of a Yang–Baxter map.
This paper is organized as follows. In section 2, as the main building block of our construction, we introduce a discrete Lax matrix, which is motivated by a binary Bäcklund–Darboux transformation [16, 17, 18] for the continuous derivative nonlinear Schrödinger (Chen–Lee–Liu [19]) hierarchy. By solving the discrete zero-curvature condition based on this Lax matrix, we obtain a fully discrete analog of the massive Thirring model in light-cone (or characteristic) coordinates. We discuss how this discrete massive Thirring model reduces to the continuous massive Thirring model in a suitable continuous limit. In section 3, we present a binary Bäcklund–Darboux transformation for the discrete massive Thirring model, which can be used to construct its explicit solutions. In section 4, we present some symmetry properties of our discrete Lax matrix and, using these properties, we translate the results in section 2 to the form of a Yang–Baxter map that admits a continuous limit. Section 5 is devoted to concluding remarks.
2 Massive Thirring model
2.1 Continuous model
The Lax-pair representation for the continuous massive Thirring model in light-cone coordinates [20, 21, 22] (or its matrix generalization [23, 24]) is given by
$\mathrm{i}\left[\begin{array}[]{c}\Psi_{1}\\ \Psi_{2}\\ \end{array}\right]_{x}=\left[\begin{array}[]{cc}-qr&\zeta q\\ \zeta r&-\zeta^{2}I\\ \end{array}\right]\left[\begin{array}[]{c}\Psi_{1}\\ \Psi_{2}\\ \end{array}\right],$ (2.7)
and
$\mathrm{i}\left[\begin{array}[]{c}\Psi_{1}\\ \Psi_{2}\\ \end{array}\right]_{t}=\left[\begin{array}[]{cc}-\frac{1}{\zeta^{2}}I&\frac{1}{\zeta}u\\[2.84526pt] \frac{1}{\zeta}v&-vu\\ \end{array}\right]\left[\begin{array}[]{c}\Psi_{1}\\ \Psi_{2}\\ \end{array}\right].$ (2.14)
Here, $\zeta$ is a constant spectral parameter and the subscripts $x$ and $t$ denote the partial differentiation with respect to these variables. The Lax-pair representation for the massive Thirring model in laboratory coordinates [25] can be obtained by applying a linear transformation of the independent variables. In this paper, unless otherwise specified, every square matrix is partitioned into four blocks as an $(M+N)\times(M+N)$ matrix; we omit the subscripts of the identity matrix $I$ and the zero matrix $O$ to denote their matrix dimensions.
The compatibility condition for the overdetermined linear equations (2.7) and (2.14) provides the equations of motion for the massive Thirring model in light-cone coordinates [20, 21, 22] (see [23, 24] for the matrix generalization):
$\left\{\begin{split}&\mathrm{i}q_{t}+u-qvu=O,\\[1.42262pt] &\mathrm{i}r_{t}-v+vur=O,\\[1.42262pt] &\mathrm{i}u_{x}-q+qru=O,\\[1.42262pt] &\mathrm{i}v_{x}+r-vqr=O.\end{split}\right.$ (2.15)
Here and hereafter, we use the symbol italic $O$, instead of 00, on the right-hand sides of the equations to imply that the dependent variables appearing on the left-hand sides can be matrix-valued.
In the case of the scalar dependent variables, by imposing the complex conjugation reduction
$r=q^{\ast},\hskip 14.22636ptv=u^{\ast},$
where the asterisk denotes the complex conjugate, we obtain a familiar form of the massive Thirring model in light-cone coordinates:
$\left\{\begin{split}&\mathrm{i}q_{t}+u-|u|^{2}q=0,\\[1.42262pt] &\mathrm{i}u_{x}-q+|q|^{2}u=0.\end{split}\right.$ (2.16)
In the case of the matrix dependent variables, by imposing the Hermitian conjugation reduction
$r=q^{\dagger},\hskip 14.22636ptv=u^{\dagger},$
where the dagger denotes the Hermitian conjugate, we obtain the matrix generalization of (2.16):
$\left\{\begin{split}&\mathrm{i}q_{t}+u-qu^{\dagger}u=O,\\[1.42262pt] &\mathrm{i}u_{x}-q+qq^{\dagger}u=O.\end{split}\right.$ (2.17)
The spectral problem (2.7) is associated with the continuous derivative nonlinear Schrödinger (known as the Chen–Lee–Liu [19]) hierarchy [26, 27, 28] (see [23, 24] for the matrix case); the massive Thirring model in light-cone coordinates can be considered as the first negative flow of the continuous Chen–Lee–Liu hierarchy [29, 23].
In the remaining part of this section, we consider how the above Lax-pair representation for the massive Thirring model in light-cone coordinates can be fully discretized.
2.2 Discrete Lax pair
A proper discretization of the Chen–Lee–Liu spectral problem (2.7) can be obtained from a suitable Bäcklund–Darboux transformation for the continuous Chen–Lee–Liu hierarchy [30]. Specifically, we consider the following discrete spectral problem:
$\left[\begin{array}[]{c}\Psi_{1,n+1}\\ \Psi_{2,n+1}\\ \end{array}\right]={\cal L}(q_{n},r_{n};\mu,\nu,\xi,\eta)\left[\begin{array}[]{c}\Psi_{1,n}\\ \Psi_{2,n}\\ \end{array}\right],$ (2.18)
where $n\in\mathbb{Z}$ represents the discrete spatial coordinate and the discrete spatial Lax matrix ${\cal L}(q_{n},r_{n};\mu,\nu,\xi,\eta)$ is defined as
${\cal L}(q_{n},r_{n};\mu,\nu,\xi,\eta)$
$:=\left[\begin{array}[]{cc}\mu I&O\\ O&\xi I\\ \end{array}\right]+\frac{\mu\nu-\xi\eta}{\xi\zeta^{2}+\nu}\left[\begin{array}[]{cc}q_{n}r_{n}(\xi\eta I-q_{n}r_{n})^{-1}&-\xi\zeta q_{n}(\xi\eta I-r_{n}q_{n})^{-1}\\ -\xi\zeta r_{n}(\xi\eta I-q_{n}r_{n})^{-1}&\xi^{2}\zeta^{2}(\xi\eta I-r_{n}q_{n})^{-1}\\ \end{array}\right]$ (2.23)
$\hphantom{:}=\left[\begin{array}[]{cc}\mu I&O\\ O&\xi I\\ \end{array}\right]+\frac{\mu\nu-\xi\eta}{\xi\zeta^{2}+\nu}\left[\begin{array}[]{c}q_{n}\\ -\xi\zeta I\\ \end{array}\right](\xi\eta I-r_{n}q_{n})^{-1}\left[\begin{array}[]{cc}r_{n}&-\xi\zeta I\\ \end{array}\right].$ (2.29)
Here, $\zeta$ is the spectral parameter common to all the Lax matrices appearing in this paper, so the dependence of ${\cal L}$ on $\zeta$ is suppressed; $\mu$, $\nu$, $\xi$ and $\eta$ are nonzero constant parameters satisfying the condition $\mu\nu-\xi\eta\neq 0$. The matrix (2.29) is equivalent, up to an unessential overall factor, to the spatial Lax matrix considered in our previous paper [30].
We use the tilde $\widetilde{}$ to denote the forward shift ($m\to m+1$) in the discrete time coordinate $m\in{\mathbb{Z}}$ [31] and consider the following discrete-time evolution:
$\left[\begin{array}[]{c}\widetilde{\Psi}_{1,n}\\ \widetilde{\Psi}_{2,n}\\ \end{array}\right]={\cal L}(u_{n},v_{n};\alpha,\beta,\gamma,\delta)\left[\begin{array}[]{c}\Psi_{1,n}\\ \Psi_{2,n}\\ \end{array}\right],$ (2.30)
where the discrete temporal Lax matrix ${\cal L}(u_{n},v_{n};\alpha,\beta,\gamma,\delta)$ is defined exactly in the same manner as in (2.29):
${\cal L}(u_{n},v_{n};\alpha,\beta,\gamma,\delta)$
$:=\left[\begin{array}[]{cc}\alpha I&O\\ O&\gamma I\\ \end{array}\right]+\frac{\alpha\beta-\gamma\delta}{\gamma\zeta^{2}+\beta}\left[\begin{array}[]{cc}u_{n}v_{n}(\gamma\delta I-u_{n}v_{n})^{-1}&-\gamma\zeta u_{n}(\gamma\delta I-v_{n}u_{n})^{-1}\\ -\gamma\zeta v_{n}(\gamma\delta I-u_{n}v_{n})^{-1}&\gamma^{2}\zeta^{2}(\gamma\delta I-v_{n}u_{n})^{-1}\\ \end{array}\right]$ (2.35)
$\hphantom{:}=\left[\begin{array}[]{cc}\alpha I&O\\ O&\gamma I\\ \end{array}\right]+\frac{\alpha\beta-\gamma\delta}{\gamma\zeta^{2}+\beta}\left[\begin{array}[]{c}u_{n}\\ -\gamma\zeta I\\ \end{array}\right](\gamma\delta I-v_{n}u_{n})^{-1}\left[\begin{array}[]{cc}v_{n}&-\gamma\zeta I\\ \end{array}\right].$ (2.41)
Here, $\alpha$, $\beta$, $\gamma$ and $\delta$ are nonzero constant parameters satisfying the condition $\alpha\beta-\gamma\delta\neq 0$.
The compatibility condition for the overdetermined linear equations (2.18) and (2.30) is given by [33, 34, 32]
${\cal L}(\widetilde{q}_{n},\widetilde{r}_{n};\mu,\nu,\xi,\eta){\cal L}(u_{n},v_{n};\alpha,\beta,\gamma,\delta)={\cal L}(u_{n+1},v_{n+1};\alpha,\beta,\gamma,\delta){\cal L}(q_{n},r_{n};\mu,\nu,\xi,\eta),$ (2.42)
which is a fully discrete analog of the zero-curvature condition. Hereafter, we assume that the nonzero parameters $\mu$, $\nu$, $\xi$, $\eta$, $\alpha$, $\beta$, $\gamma$ and $\delta$ satisfy the conditions $\alpha\nu-\delta\xi\neq 0$, $\alpha\eta-\delta\mu\neq 0$, $\beta\mu-\gamma\eta\neq 0$ and $\beta\xi-\gamma\nu\neq 0$, in addition to the conditions $\mu\nu-\xi\eta\neq 0$ and $\alpha\beta-\gamma\delta\neq 0$. In a nutshell, the values of $\nu/\xi$, $\eta/\mu$, $\beta/\gamma$ and $\delta/\alpha$ are assumed to be all different.
${\cal L}(q_{n},r_{n};\mu,\nu,\xi,\eta)$${\cal L}(\widetilde{q}_{n},\widetilde{r}_{n};\mu,\nu,\xi,\eta)$${\cal L}(u_{n},v_{n};\alpha,\beta,\gamma,\delta)$${\cal L}(u_{n+1},v_{n+1};\alpha,\beta,\gamma,\delta)$$\left[\begin{array}[]{c}\Psi_{1,n}\\ \Psi_{2,n}\\ \end{array}\right]$$\left[\begin{array}[]{c}\widetilde{\Psi}_{1,n}\\ \widetilde{\Psi}_{2,n}\\ \end{array}\right]$$\left[\begin{array}[]{c}\Psi_{1,n+1}\\ \Psi_{2,n+1}\\ \end{array}\right]$$\left[\begin{array}[]{c}\widetilde{\Psi}_{1,n+1}\\ \widetilde{\Psi}_{2,n+1}\\ \end{array}\right]$
Clearly, we can multiply the discrete spatial/temporal Lax matrix by any nonzero constant factor without changing the discrete zero-curvature condition (2.42). Alternatively but equivalently, we can consider the gauge transformation:
$\left[\begin{array}[]{c}\Psi_{1,n}\\ \Psi_{2,n}\\ \end{array}\right]\to a^{n}b^{m}\left[\begin{array}[]{c}\Psi_{1,n}\\ \Psi_{2,n}\\ \end{array}\right],$
where $m$ is the discrete time coordinate, and $a$ and $b$ are arbitrary nonzero constant factors that may depend on the spectral parameter $\zeta$. It is often more convenient to adopt the temporal Lax matrix of the following rescaled form:
$\frac{\zeta+\frac{\beta}{\gamma\zeta}}{\frac{\alpha}{\delta}\zeta+\frac{1}{\zeta}}{\cal L}(u_{n},v_{n};\alpha,\beta,\gamma,\delta)$
$=\left[\begin{array}[]{cc}\delta I&O\\ O&\beta I\\ \end{array}\right]+\frac{\alpha\beta-\gamma\delta}{\alpha+\frac{\delta}{\zeta^{2}}}\left[\begin{array}[]{cc}\frac{\delta^{2}}{\zeta^{2}}(\gamma\delta I-u_{n}v_{n})^{-1}&-\frac{\delta}{\zeta}u_{n}(\gamma\delta I-v_{n}u_{n})^{-1}\\[2.84526pt] -\frac{\delta}{\zeta}v_{n}(\gamma\delta I-u_{n}v_{n})^{-1}&v_{n}u_{n}(\gamma\delta I-v_{n}u_{n})^{-1}\\ \end{array}\right]$ (2.47)
$=\left[\begin{array}[]{cc}\delta I&O\\ O&\beta I\\ \end{array}\right]+\frac{\alpha\beta-\gamma\delta}{\alpha+\frac{\delta}{\zeta^{2}}}\left[\begin{array}[]{c}-\frac{\delta}{\zeta}I\\ v_{n}\\ \end{array}\right](\gamma\delta I-u_{n}v_{n})^{-1}\left[\begin{array}[]{cc}-\frac{\delta}{\zeta}I&u_{n}\\ \end{array}\right].$ (2.53)
We note that
[…]
$=I+\varDelta\frac{\nu-\eta}{\varDelta\zeta^{2}+\nu}\left[\begin{array}[]{cc}q_{n}r_{n}(\eta I-\varDelta q_{n}r_{n})^{-1}&-\zeta q_{n}(\eta I-\varDelta r_{n}q_{n})^{-1}\\ -\zeta r_{n}(\eta I-\varDelta q_{n}r_{n})^{-1}&\zeta^{2}(\eta I-\varDelta r_{n}q_{n})^{-1}\\ \end{array}\right]$ (2.56)
$=I+\varDelta\frac{\nu-\eta}{\nu\eta}\left[\begin{array}[]{cc}q_{n}r_{n}&-\zeta q_{n}\\ -\zeta r_{n}&\zeta^{2}I\\ \end{array}\right]+\mathcal{O}(\varDelta^{2}),$ (2.59)
and
$\hphantom{=}\frac{\zeta+\frac{h}{\gamma\zeta}}{\frac{\alpha}{h}\zeta+\frac{1}{\zeta}}{\cal L}(u_{n},v_{n};\alpha/h,1,\gamma/h,1)$
$=I+h\frac{\alpha-\gamma}{\alpha+\frac{h}{\zeta^{2}}}\left[\begin{array}[]{cc}\frac{1}{\zeta^{2}}(\gamma I-hu_{n}v_{n})^{-1}&-\frac{1}{\zeta}u_{n}(\gamma I-hv_{n}u_{n})^{-1}\\[2.84526pt] -\frac{1}{\zeta}v_{n}(\gamma I-hu_{n}v_{n})^{-1}&v_{n}u_{n}(\gamma I-hv_{n}u_{n})^{-1}\\ \end{array}\right]$ (2.62)
$=I+h\frac{\alpha-\gamma}{\alpha\gamma}\left[\begin{array}[]{cc}\frac{1}{\zeta^{2}}I&-\frac{1}{\zeta}u_{n}\\[2.84526pt] -\frac{1}{\zeta}v_{n}&v_{n}u_{n}\\ \end{array}\right]+\mathcal{O}(h^{2}).$ (2.65)
Thus, we can recover the spatial Lax matrix in (2.7) by setting
$\frac{\nu-\eta}{\nu\eta}=\mathrm{i},$
and taking the continuous space limit $\varDelta\to 0$; the temporal Lax matrix in (2.14) can be recovered by setting
$\frac{\alpha-\gamma}{\alpha\gamma}=\mathrm{i},$
and taking the continuous time limit $h\to 0$.
2.3 Fully discrete equations of motion
Substituting (2.29) and (2.41) into the discrete zero-curvature condition (2.42), we obtain
$\left\{\left[\begin{array}[]{cc}\mu I&O\\ O&\xi I\\ \end{array}\right]+\frac{\mu\nu-\xi\eta}{\xi\zeta^{2}+\nu}\left[\begin{array}[]{c}\widetilde{q}_{n}\\ -\xi\zeta I\\ \end{array}\right](\xi\eta I-\widetilde{r}_{n}\widetilde{q}_{n})^{-1}\left[\begin{array}[]{cc}\widetilde{r}_{n}&-\xi\zeta I\\ \end{array}\right]\right\}$ (2.71)
×$\times\left\{\left[\begin{array}[]{cc}\alpha I&O\\ O&\gamma I\\ \end{array}\right]+\frac{\alpha\beta-\gamma\delta}{\gamma\zeta^{2}+\beta}\left[\begin{array}[]{c}u_{n}\\ -\gamma\zeta I\\ \end{array}\right](\gamma\delta I-v_{n}u_{n})^{-1}\left[\begin{array}[]{cc}v_{n}&-\gamma\zeta I\\ \end{array}\right]\right\}$ (2.77)
$=$ $\left\{\left[\begin{array}[]{cc}\alpha I&O\\ O&\gamma I\\ \end{array}\right]+\frac{\alpha\beta-\gamma\delta}{\gamma\zeta^{2}+\beta}\left[\begin{array}[]{c}u_{n+1}\\ -\gamma\zeta I\\ \end{array}\right](\gamma\delta I-v_{n+1}u_{n+1})^{-1}\left[\begin{array}[]{cc}v_{n+1}&-\gamma\zeta I\\ \end{array}\right]\right\}$ (2.83)
×$\times\left\{\left[\begin{array}[]{cc}\mu I&O\\ O&\xi I\\ \end{array}\right]+\frac{\mu\nu-\xi\eta}{\xi\zeta^{2}+\nu}\left[\begin{array}[]{c}q_{n}\\ -\xi\zeta I\\ \end{array}\right](\xi\eta I-r_{n}q_{n})^{-1}\left[\begin{array}[]{cc}r_{n}&-\xi\zeta I\\ \end{array}\right]\right\}.$ (2.89)
We subtract the common term:
$\left[\begin{array}[]{cc}\alpha\mu I&O\\ O&\gamma\xi I\\ \end{array}\right]$
from both sides of (2.89) and multiply it both from the left and right by the diagonal matrix:
$\left[\begin{array}[]{cc}I&O\\ O&\frac{1}{\zeta}I\\ \end{array}\right].$
Then, because of the condition $\beta\xi-\gamma\nu\neq 0$ (or, equivalently, $\nu/\xi\neq\beta/\gamma$) and the identities:
$\frac{1}{(\xi\zeta^{2}+\nu)(\gamma\zeta^{2}+\beta)}=\frac{1}{\beta\xi-\gamma\nu}\left(\frac{\xi}{\xi\zeta^{2}+\nu}-\frac{\gamma}{\gamma\zeta^{2}+\beta}\right),$
$\frac{\zeta^{2}}{(\xi\zeta^{2}+\nu)(\gamma\zeta^{2}+\beta)}=\frac{1}{\beta\xi-\gamma\nu}\left(-\frac{\nu}{\xi\zeta^{2}+\nu}+\frac{\beta}{\gamma\zeta^{2}+\beta}\right),$
we obtain from the coefficients of $1/(\xi\zeta^{2}+\nu)$ and $1/(\gamma\zeta^{2}+\beta)$ the following two relations:
$\left[\begin{array}[]{c}\widetilde{q}_{n}\\ -\xi I\\ \end{array}\right](\xi\eta I-\widetilde{r}_{n}\widetilde{q}_{n})^{-1}\left[\begin{array}[]{cc}\widetilde{r}_{n}&-\xi I\\ \end{array}\right]$ (2.93)
×$\times\left\{\left[\begin{array}[]{cc}\alpha I&O\\ O&\gamma I\\ \end{array}\right]+\frac{\alpha\beta-\gamma\delta}{\beta\xi-\gamma\nu}\left[\begin{array}[]{cc}\xi I&O\\ O&-\nu I\\ \end{array}\right]\left[\begin{array}[]{c}u_{n}\\ -\gamma I\\ \end{array}\right](\gamma\delta I-v_{n}u_{n})^{-1}\left[\begin{array}[]{cc}v_{n}&-\gamma I\\ \end{array}\right]\right\}$ (2.101)
=$=\left\{\left[\begin{array}[]{cc}\alpha I&O\\ O&\gamma I\\ \end{array}\right]+\frac{\alpha\beta-\gamma\delta}{\beta\xi-\gamma\nu}\left[\begin{array}[]{c}u_{n+1}\\ -\gamma I\\ \end{array}\right](\gamma\delta I-v_{n+1}u_{n+1})^{-1}\left[\begin{array}[]{cc}v_{n+1}&-\gamma I\\ \end{array}\right]\left[\begin{array}[]{cc}\xi I&O\\ O&-\nu I\\ \end{array}\right]\right\}$ (2.109)
×$\times\left[\begin{array}[]{c}q_{n}\\ -\xi I\\ \end{array}\right](\xi\eta I-r_{n}q_{n})^{-1}\left[\begin{array}[]{cc}r_{n}&-\xi I\\ \end{array}\right],$ (2.113)
and
$\left\{\left[\begin{array}[]{cc}\mu I&O\\ O&\xi I\\ \end{array}\right]+\frac{\mu\nu-\xi\eta}{\beta\xi-\gamma\nu}\left[\begin{array}[]{c}\widetilde{q}_{n}\\ -\xi I\\ \end{array}\right](\xi\eta I-\widetilde{r}_{n}\widetilde{q}_{n})^{-1}\left[\begin{array}[]{cc}\widetilde{r}_{n}&-\xi I\\ \end{array}\right]\left[\begin{array}[]{cc}-\gamma I&O\\ O&\beta I\\ \end{array}\right]\right\}$ (2.121)
×$\times\left[\begin{array}[]{c}u_{n}\\ -\gamma I\\ \end{array}\right](\gamma\delta I-v_{n}u_{n})^{-1}\left[\begin{array}[]{cc}v_{n}&-\gamma I\\ \end{array}\right]$ (2.125)
$=\left[\begin{array}[]{c}u_{n+1}\\ -\gamma I\\ \end{array}\right](\gamma\delta I-v_{n+1}u_{n+1})^{-1}\left[\begin{array}[]{cc}v_{n+1}&-\gamma I\\ \end{array}\right]$ (2.129)
×$\times\left\{\left[\begin{array}[]{cc}\mu I&O\\ O&\xi I\\ \end{array}\right]+\frac{\mu\nu-\xi\eta}{\beta\xi-\gamma\nu}\left[\begin{array}[]{cc}-\gamma I&O\\ O&\beta I\\ \end{array}\right]\left[\begin{array}[]{c}q_{n}\\ -\xi I\\ \end{array}\right](\xi\eta I-r_{n}q_{n})^{-1}\left[\begin{array}[]{cc}r_{n}&-\xi I\\ \end{array}\right]\right\}.$ (2.137)
Noting the identities:
$\left\{\left[\begin{array}[]{cc}\alpha I&O\\ O&\gamma I\\ \end{array}\right]+\frac{\alpha\beta-\gamma\delta}{\beta\xi-\gamma\nu}\left[\begin{array}[]{cc}\xi I&O\\ O&-\nu I\\ \end{array}\right]\left[\begin{array}[]{c}u_{n}\\ -\gamma I\\ \end{array}\right](\gamma\delta I-v_{n}u_{n})^{-1}\left[\begin{array}[]{cc}v_{n}&-\gamma I\\ \end{array}\right]\right\}$ (2.145)
×$\times\left\{\left[\begin{array}[]{cc}\gamma I&O\\ O&\alpha I\\ \end{array}\right]+\frac{\alpha\beta-\gamma\delta}{\alpha\nu-\delta\xi}\left[\begin{array}[]{cc}\xi I&O\\ O&-\nu I\\ \end{array}\right]\left[\begin{array}[]{c}u_{n}\\ -\alpha I\\ \end{array}\right](\alpha\beta I-v_{n}u_{n})^{-1}\left[\begin{array}[]{cc}v_{n}&-\alpha I\\ \end{array}\right]\right\}$ (2.153)
$=\alpha\gamma I,$ (2.154)
and
$\left\{\left[\begin{array}[]{cc}\mu I&O\\ O&\xi I\\ \end{array}\right]+\frac{\mu\nu-\xi\eta}{\beta\xi-\gamma\nu}\left[\begin{array}[]{cc}-\gamma I&O\\ O&\beta I\\ \end{array}\right]\left[\begin{array}[]{c}q_{n}\\ -\xi I\\ \end{array}\right](\xi\eta I-r_{n}q_{n})^{-1}\left[\begin{array}[]{cc}r_{n}&-\xi I\\ \end{array}\right]\right\}$ (2.162)
×$\times\left\{\left[\begin{array}[]{cc}\xi I&O\\ O&\mu I\\ \end{array}\right]+\frac{\mu\nu-\xi\eta}{\gamma\eta-\beta\mu}\left[\begin{array}[]{cc}-\gamma I&O\\ O&\beta I\\ \end{array}\right]\left[\begin{array}[]{c}q_{n}\\ -\mu I\\ \end{array}\right](\mu\nu I-r_{n}q_{n})^{-1}\left[\begin{array}[]{cc}r_{n}&-\mu I\\ \end{array}\right]\right\}$ (2.170)
$=\mu\xi I,$ (2.171)
we can rewrite (2.113) and (2.137) as
$\left[\begin{array}[]{c}\widetilde{q}_{n}\\ -\xi I\\ \end{array}\right](\xi\eta I-\widetilde{r}_{n}\widetilde{q}_{n})^{-1}\left[\begin{array}[]{cc}\widetilde{r}_{n}&-\xi I\\ \end{array}\right]$ (2.175)
$=\frac{1}{\alpha\gamma}\left\{\left[\begin{array}[]{cc}\alpha I&O\\ O&\gamma I\\ \end{array}\right]+\frac{\alpha\beta-\gamma\delta}{\beta\xi-\gamma\nu}\left[\begin{array}[]{c}u_{n+1}\\ -\gamma I\\ \end{array}\right](\gamma\delta I-v_{n+1}u_{n+1})^{-1}\left[\begin{array}[]{cc}v_{n+1}&-\gamma I\\ \end{array}\right]\left[\begin{array}[]{cc}\xi I&O\\ O&-\nu I\\ \end{array}\right]\right\}$ (2.183)
×$\times\left[\begin{array}[]{c}q_{n}\\ -\xi I\\ \end{array}\right](\xi\eta I-r_{n}q_{n})^{-1}\left[\begin{array}[]{cc}r_{n}&-\xi I\\ \end{array}\right]$ (2.187)
×$\times\left\{\left[\begin{array}[]{cc}\gamma I&O\\ O&\alpha I\\ \end{array}\right]+\frac{\alpha\beta-\gamma\delta}{\alpha\nu-\delta\xi}\left[\begin{array}[]{cc}\xi I&O\\ O&-\nu I\\ \end{array}\right]\left[\begin{array}[]{c}u_{n}\\ -\alpha I\\ \end{array}\right](\alpha\beta I-v_{n}u_{n})^{-1}\left[\begin{array}[]{cc}v_{n}&-\alpha I\\ \end{array}\right]\right\},$ (2.195)
and
$\frac{1}{\mu\xi}\left\{\left[\begin{array}[]{cc}\mu I&O\\ O&\xi I\\ \end{array}\right]+\frac{\mu\nu-\xi\eta}{\beta\xi-\gamma\nu}\left[\begin{array}[]{c}\widetilde{q}_{n}\\ -\xi I\\ \end{array}\right](\xi\eta I-\widetilde{r}_{n}\widetilde{q}_{n})^{-1}\left[\begin{array}[]{cc}\widetilde{r}_{n}&-\xi I\\ \end{array}\right]\left[\begin{array}[]{cc}-\gamma I&O\\ O&\beta I\\ \end{array}\right]\right\}$ (2.203)
×$\times\left[\begin{array}[]{c}u_{n}\\ -\gamma I\\ \end{array}\right](\gamma\delta I-v_{n}u_{n})^{-1}\left[\begin{array}[]{cc}v_{n}&-\gamma I\\ \end{array}\right]$ (2.207)
×$\times\left\{\left[\begin{array}[]{cc}\xi I&O\\ O&\mu I\\ \end{array}\right]+\frac{\mu\nu-\xi\eta}{\gamma\eta-\beta\mu}\left[\begin{array}[]{cc}-\gamma I&O\\ O&\beta I\\ \end{array}\right]\left[\begin{array}[]{c}q_{n}\\ -\mu I\\ \end{array}\right](\mu\nu I-r_{n}q_{n})^{-1}\left[\begin{array}[]{cc}r_{n}&-\mu I\\ \end{array}\right]\right\}$ (2.215)
$=\left[\begin{array}[]{c}u_{n+1}\\ -\gamma I\\ \end{array}\right](\gamma\delta I-v_{n+1}u_{n+1})^{-1}\left[\begin{array}[]{cc}v_{n+1}&-\gamma I\\ \end{array}\right],$ (2.219)
respectively. Thus, we can set
$\left[\begin{array}[]{c}\widetilde{q}_{n}\\ -\xi I\\ \end{array}\right]$ (2.222)
=$=\left\{\left[\begin{array}[]{cc}\alpha I&O\\ O&\gamma I\\ \end{array}\right]+\frac{\alpha\beta-\gamma\delta}{\beta\xi-\gamma\nu}\left[\begin{array}[]{c}u_{n+1}\\ -\gamma I\\ \end{array}\right](\gamma\delta I-v_{n+1}u_{n+1})^{-1}\left[\begin{array}[]{cc}v_{n+1}&-\gamma I\\ \end{array}\right]\left[\begin{array}[]{cc}\xi I&O\\ O&-\nu I\\ \end{array}\right]\right\}$ (2.230)
×$\times\left[\begin{array}[]{c}q_{n}\\ -\xi I\\ \end{array}\right]X_{n},$ (2.233)
$\left[\begin{array}[]{cc}\widetilde{r}_{n}&-\xi I\\ \end{array}\right]=Y_{n}\left[\begin{array}[]{cc}r_{n}&-\xi I\\ \end{array}\right]$ (2.236)
×$\times\left\{\left[\begin{array}[]{cc}\gamma I&O\\ O&\alpha I\\ \end{array}\right]+\frac{\alpha\beta-\gamma\delta}{\alpha\nu-\delta\xi}\left[\begin{array}[]{cc}\xi I&O\\ O&-\nu I\\ \end{array}\right]\left[\begin{array}[]{c}u_{n}\\ -\alpha I\\ \end{array}\right](\alpha\beta I-v_{n}u_{n})^{-1}\left[\begin{array}[]{cc}v_{n}&-\alpha I\\ \end{array}\right]\right\},$ (2.244)
and
$\left[\begin{array}[]{c}u_{n+1}\\ -\gamma I\\ \end{array}\right]$ (2.247)
=$=\left\{\left[\begin{array}[]{cc}\mu I&O\\ O&\xi I\\ \end{array}\right]+\frac{\mu\nu-\xi\eta}{\beta\xi-\gamma\nu}\left[\begin{array}[]{c}\widetilde{q}_{n}\\ -\xi I\\ \end{array}\right](\xi\eta I-\widetilde{r}_{n}\widetilde{q}_{n})^{-1}\left[\begin{array}[]{cc}\widetilde{r}_{n}&-\xi I\\ \end{array}\right]\left[\begin{array}[]{cc}-\gamma I&O\\ O&\beta I\\ \end{array}\right]\right\}$ (2.255)
×$\times\left[\begin{array}[]{c}u_{n}\\ -\gamma I\\ \end{array}\right]Z_{n},$ (2.258)
$\left[\begin{array}[]{cc}v_{n+1}&-\gamma I\\ \end{array}\right]=W_{n}\left[\begin{array}[]{cc}v_{n}&-\gamma I\\ \end{array}\right]$ (2.261)
×$\times\left\{\left[\begin{array}[]{cc}\xi I&O\\ O&\mu I\\ \end{array}\right]+\frac{\mu\nu-\xi\eta}{\gamma\eta-\beta\mu}\left[\begin{array}[]{cc}-\gamma I&O\\ O&\beta I\\ \end{array}\right]\left[\begin{array}[]{c}q_{n}\\ -\mu I\\ \end{array}\right](\mu\nu I-r_{n}q_{n})^{-1}\left[\begin{array}[]{cc}r_{n}&-\mu I\\ \end{array}\right]\right\},$ (2.269)
where $X_{n}$, $Y_{n}$, $Z_{n}$ and $W_{n}$ are square matrices that satisfy the relations:
$X_{n}(\xi\eta I-\widetilde{r}_{n}\widetilde{q}_{n})^{-1}Y_{n}=\frac{1}{\alpha\gamma}(\xi\eta I-r_{n}q_{n})^{-1},$ (2.270)
and
$Z_{n}(\gamma\delta I-v_{n+1}u_{n+1})^{-1}W_{n}=\frac{1}{\mu\xi}(\gamma\delta I-v_{n}u_{n})^{-1}.$ (2.271)
Using (2.269) and (2.271), we can rewrite (2.233) as
$\left[\begin{array}[]{c}\widetilde{q}_{n}\\ -\xi I\\ \end{array}\right]X_{n}^{-1}=\left\{\left[\begin{array}[]{cc}\alpha I&O\\ O&\gamma I\\ \end{array}\right]+\frac{\alpha\beta-\gamma\delta}{\beta\xi-\gamma\nu}\left[\begin{array}[]{c}u_{n+1}\\ -\gamma I\\ \end{array}\right]\frac{1}{\mu\xi}Z_{n}^{-1}(\gamma\delta I-v_{n}u_{n})^{-1}\left[\begin{array}[]{cc}v_{n}&-\gamma I\\ \end{array}\right]\right.$ (2.279)
×$\left.\times\left[\left[\begin{array}[]{cc}\xi I&O\\ O&\mu I\\ \end{array}\right]+\frac{\mu\nu-\xi\eta}{\gamma\eta-\beta\mu}\left[\begin{array}[]{cc}-\gamma I&O\\ O&\beta I\\ \end{array}\right]\left[\begin{array}[]{c}q_{n}\\ -\mu I\\ \end{array}\right](\mu\nu I-r_{n}q_{n})^{-1}\left[\begin{array}[]{cc}r_{n}&-\mu I\\ \end{array}\right]\right]\left[\begin{array}[]{cc}\xi I&O\\ O&-\nu I\\ \end{array}\right]\right\}$ (2.289)
×$\times\left[\begin{array}[]{c}q_{n}\\ -\xi I\\ \end{array}\right]$ (2.292)
$=\left[\begin{array}[]{c}\alpha q_{n}\\ -\gamma\xi I\\ \end{array}\right]-\frac{\alpha\beta-\gamma\delta}{\beta\mu-\gamma\eta}\left[\begin{array}[]{c}u_{n+1}\\ -\gamma I\\ \end{array}\right]Z_{n}^{-1}(\gamma\delta I-v_{n}u_{n})^{-1}(\gamma\eta I-v_{n}q_{n}).$ (2.297)
Similarly, using (2.244) and (2.270), we can rewrite (2.258) as
$\left[\begin{array}[]{c}u_{n+1}\\ -\gamma I\\ \end{array}\right]Z_{n}^{-1}=\left\{\left[\begin{array}[]{cc}\mu I&O\\ O&\xi I\\ \end{array}\right]+\frac{\mu\nu-\xi\eta}{\beta\xi-\gamma\nu}\left[\begin{array}[]{c}\widetilde{q}_{n}\\ -\xi I\\ \end{array}\right]\frac{1}{\alpha\gamma}X_{n}^{-1}(\xi\eta I-r_{n}q_{n})^{-1}\left[\begin{array}[]{cc}r_{n}&-\xi I\\ \end{array}\right]\right.$ (2.305)
×$\left.\times\left[\left[\begin{array}[]{cc}\gamma I&O\\ O&\alpha I\\ \end{array}\right]+\frac{\alpha\beta-\gamma\delta}{\alpha\nu-\delta\xi}\left[\begin{array}[]{cc}\xi I&O\\ O&-\nu I\\ \end{array}\right]\left[\begin{array}[]{c}u_{n}\\ -\alpha I\\ \end{array}\right](\alpha\beta I-v_{n}u_{n})^{-1}\left[\begin{array}[]{cc}v_{n}&-\alpha I\\ \end{array}\right]\right]\left[\begin{array}[]{cc}-\gamma I&O\\ O&\beta I\\ \end{array}\right]\right\}$ (2.315)
×$\times\left[\begin{array}[]{c}u_{n}\\ -\gamma I\\ \end{array}\right]$ (2.318)
$=\left[\begin{array}[]{c}\mu u_{n}\\ -\gamma\xi I\\ \end{array}\right]-\frac{\mu\nu-\xi\eta}{\alpha\nu-\delta\xi}\left[\begin{array}[]{c}\widetilde{q}_{n}\\ -\xi I\\ \end{array}\right]X_{n}^{-1}(\xi\eta I-r_{n}q_{n})^{-1}(\delta\xi I-r_{n}u_{n}).$ (2.323)
Combining (2.297) with (2.323), we obtain
$\left[\begin{array}[]{c}\widetilde{q}_{n}\\ -\xi I\\ \end{array}\right]X_{n}^{-1}\left\{I-\frac{(\alpha\beta-\gamma\delta)(\mu\nu-\xi\eta)}{(\beta\mu-\gamma\eta)(\alpha\nu-\delta\xi)}(\xi\eta I-r_{n}q_{n})^{-1}(\delta\xi I-r_{n}u_{n})(\gamma\delta I-v_{n}u_{n})^{-1}(\gamma\eta I-v_{n}q_{n})\right\}$ (2.326)
$=\left[\begin{array}[]{c}\alpha q_{n}\\ -\gamma\xi I\\ \end{array}\right]-\frac{\alpha\beta-\gamma\delta}{\beta\mu-\gamma\eta}\left[\begin{array}[]{c}\mu u_{n}\\ -\gamma\xi I\\ \end{array}\right](\gamma\delta I-v_{n}u_{n})^{-1}(\gamma\eta I-v_{n}q_{n}),$ (2.331)
and
$\left[\begin{array}[]{c}u_{n+1}\\ -\gamma I\\ \end{array}\right]Z_{n}^{-1}\left\{I-\frac{(\alpha\beta-\gamma\delta)(\mu\nu-\xi\eta)}{(\beta\mu-\gamma\eta)(\alpha\nu-\delta\xi)}(\gamma\delta I-v_{n}u_{n})^{-1}(\gamma\eta I-v_{n}q_{n})(\xi\eta I-r_{n}q_{n})^{-1}(\delta\xi I-r_{n}u_{n})\right\}$ (2.334)
$=\left[\begin{array}[]{c}\mu u_{n}\\ -\gamma\xi I\\ \end{array}\right]-\frac{\mu\nu-\xi\eta}{\alpha\nu-\delta\xi}\left[\begin{array}[]{c}\alpha q_{n}\\ -\gamma\xi I\\ \end{array}\right](\xi\eta I-r_{n}q_{n})^{-1}(\delta\xi I-r_{n}u_{n}).$ (2.339)
Equations (2.331) and (2.339) imply
$\widetilde{q}_{n}=$ $\;(\gamma\delta I-u_{n}v_{n})^{-1}\left[-\alpha\delta(\beta\mu-\gamma\eta)q_{n}+\mu\eta(\alpha\beta-\gamma\delta)u_{n}-(\alpha\eta-\delta\mu)u_{n}v_{n}q_{n}\right]$
×$\times\left[\beta\gamma(\alpha\eta-\delta\mu)I+(\beta\mu-\gamma\eta)v_{n}u_{n}-(\alpha\beta-\gamma\delta)v_{n}q_{n}\right]^{-1}(\gamma\delta I-v_{n}u_{n}),$ (2.340)
and
$u_{n+1}=$ $\;(\xi\eta I-q_{n}r_{n})^{-1}\left[\mu\eta(\alpha\nu-\delta\xi)u_{n}-\alpha\delta(\mu\nu-\xi\eta)q_{n}-(\alpha\eta-\delta\mu)q_{n}r_{n}u_{n}\right]$
×$\times\left[\nu\xi(\alpha\eta-\delta\mu)I-(\alpha\nu-\delta\xi)r_{n}q_{n}+(\mu\nu-\xi\eta)r_{n}u_{n}\right]^{-1}(\xi\eta I-r_{n}q_{n}),$ (2.341)
respectively, while equations (2.244) and (2.269) imply
$\widetilde{r}_{n}=$ $\;(\alpha\beta I-v_{n}u_{n})\left[\alpha\delta(\beta\xi-\gamma\nu)I+(\alpha\nu-\delta\xi)v_{n}u_{n}-(\alpha\beta-\gamma\delta)r_{n}u_{n}\right]^{-1}$
×$\times\left[-\beta\gamma(\alpha\nu-\delta\xi)r_{n}+\nu\xi(\alpha\beta-\gamma\delta)v_{n}-(\beta\xi-\gamma\nu)r_{n}u_{n}v_{n}\right](\alpha\beta I-u_{n}v_{n})^{-1},$ (2.342)
and
$v_{n+1}=$ $\;(\mu\nu I-r_{n}q_{n})\left[\mu\eta(\beta\xi-\gamma\nu)I-(\beta\mu-\gamma\eta)r_{n}q_{n}+(\mu\nu-\xi\eta)v_{n}q_{n}\right]^{-1}$
×$\times\left[\nu\xi(\beta\mu-\gamma\eta)v_{n}-\beta\gamma(\mu\nu-\xi\eta)r_{n}-(\beta\xi-\gamma\nu)v_{n}q_{n}r_{n}\right](\mu\nu I-q_{n}r_{n})^{-1},$ (2.343)
respectively.
With a lengthy (but almost straightforward) calculation, we can show that relation (2.270) is a direct consequence of (2.244) and (2.331), i.e.,
$Y_{n}^{-1}(\widetilde{r}_{n}\widetilde{q}_{n}-\xi\eta I)X_{n}^{-1}(\xi\eta I-r_{n}q_{n})^{-1}$
$=Y_{n}^{-1}\left[\begin{array}[]{cc}\widetilde{r}_{n}&-\xi I\\ \end{array}\right]\left[\begin{array}[]{cc}I&O\\ O&-\frac{\eta}{\xi}I\\ \end{array}\right]\left[\begin{array}[]{c}\widetilde{q}_{n}\\ -\xi I\\ \end{array}\right]X_{n}^{-1}(\xi\eta I-r_{n}q_{n})^{-1}$ (2.349)
$=\left[\begin{array}[]{cc}r_{n}&-\xi I\\ \end{array}\right]\left\{\left[\begin{array}[]{cc}\gamma I&O\\ O&\alpha I\\ \end{array}\right]+\frac{\alpha\beta-\gamma\delta}{\alpha\nu-\delta\xi}\left[\begin{array}[]{cc}\xi I&O\\ O&-\nu I\\ \end{array}\right]\left[\begin{array}[]{c}u_{n}\\ -\alpha I\\ \end{array}\right](\alpha\beta I-v_{n}u_{n})^{-1}\left[\begin{array}[]{cc}v_{n}&-\alpha I\\ \end{array}\right]\right\}$ (2.358)
×$\times\left\{\left[\begin{array}[]{c}\alpha q_{n}\\ \gamma\eta I\\ \end{array}\right]-\frac{\alpha\beta-\gamma\delta}{\beta\mu-\gamma\eta}\left[\begin{array}[]{c}\mu u_{n}\\ \gamma\eta I\\ \end{array}\right](\gamma\delta I-v_{n}u_{n})^{-1}(\gamma\eta I-v_{n}q_{n})\right\}$ (2.363)
×$\times\left\{\xi\eta I-r_{n}q_{n}-\frac{(\alpha\beta-\gamma\delta)(\mu\nu-\xi\eta)}{(\beta\mu-\gamma\eta)(\alpha\nu-\delta\xi)}(\delta\xi I-r_{n}u_{n})(\gamma\delta I-v_{n}u_{n})^{-1}(\gamma\eta I-v_{n}q_{n})\right\}^{-1}$
$=-\alpha\gamma I.$
Similarly, we can show that relation (2.271) is a direct consequence of (2.269) and (2.339), i.e.,
$W_{n}^{-1}(v_{n+1}u_{n+1}-\gamma\delta I)Z_{n}^{-1}(\gamma\delta I-v_{n}u_{n})^{-1}$
$=W_{n}^{-1}\left[\begin{array}[]{cc}v_{n+1}&-\gamma I\\ \end{array}\right]\left[\begin{array}[]{cc}I&O\\ O&-\frac{\delta}{\gamma}I\\ \end{array}\right]\left[\begin{array}[]{c}u_{n+1}\\ -\gamma I\\ \end{array}\right]Z_{n}^{-1}(\gamma\delta I-v_{n}u_{n})^{-1}$ (2.369)
$=\left[\begin{array}[]{cc}v_{n}&-\gamma I\\ \end{array}\right]\left\{\left[\begin{array}[]{cc}\xi I&O\\ O&\mu I\\ \end{array}\right]+\frac{\mu\nu-\xi\eta}{\gamma\eta-\beta\mu}\left[\begin{array}[]{cc}-\gamma I&O\\ O&\beta I\\ \end{array}\right]\left[\begin{array}[]{c}q_{n}\\ -\mu I\\ \end{array}\right](\mu\nu I-r_{n}q_{n})^{-1}\left[\begin{array}[]{cc}r_{n}&-\mu I\\ \end{array}\right]\right\}$ (2.378)
×$\times\left\{\left[\begin{array}[]{c}\mu u_{n}\\ \delta\xi I\\ \end{array}\right]-\frac{\mu\nu-\xi\eta}{\alpha\nu-\delta\xi}\left[\begin{array}[]{c}\alpha q_{n}\\ \delta\xi I\\ \end{array}\right](\xi\eta I-r_{n}q_{n})^{-1}(\delta\xi I-r_{n}u_{n})\right\}$ (2.383)
×$\times\left\{\gamma\delta I-v_{n}u_{n}-\frac{(\alpha\beta-\gamma\delta)(\mu\nu-\xi\eta)}{(\beta\mu-\gamma\eta)(\alpha\nu-\delta\xi)}(\gamma\eta I-v_{n}q_{n})(\xi\eta I-r_{n}q_{n})^{-1}(\delta\xi I-r_{n}u_{n})\right\}^{-1}$
$=-\mu\xi I.$
To summarize, the discrete zero-curvature condition (2.42) for the Lax pair given by (2.29) and (2.41) is equivalent to the fully discrete system of equations (2.340)–(2.343).
2.4 Complex conjugation reduction and continuous limit
Note that (2.342) and (2.343) admit the alternative expressions:
$\widetilde{r}_{n}=$ $\;(\gamma\delta I-v_{n}u_{n})^{-1}\left[-\beta\gamma(\alpha\nu-\delta\xi)r_{n}+\nu\xi(\alpha\beta-\gamma\delta)v_{n}-(\beta\xi-\gamma\nu)v_{n}u_{n}r_{n}\right]$
×$\times\left[\alpha\delta(\beta\xi-\gamma\nu)I+(\alpha\nu-\delta\xi)u_{n}v_{n}-(\alpha\beta-\gamma\delta)u_{n}r_{n}\right]^{-1}(\gamma\delta I-u_{n}v_{n}),$ (2.384)
and
$v_{n+1}=$ $\;(\xi\eta I-r_{n}q_{n})^{-1}\left[\nu\xi(\beta\mu-\gamma\eta)v_{n}-\beta\gamma(\mu\nu-\xi\eta)r_{n}-(\beta\xi-\gamma\nu)r_{n}q_{n}v_{n}\right]$
×$\times\left[\mu\eta(\beta\xi-\gamma\nu)I-(\beta\mu-\gamma\eta)q_{n}r_{n}+(\mu\nu-\xi\eta)q_{n}v_{n}\right]^{-1}(\xi\eta I-q_{n}r_{n}),$ (2.385)
respectively. In fact, (2.384) can be obtained from (2.244) with (2.154), while (2.385) can be obtained from (2.269) with (2.171).
In the case of $\beta=\alpha^{\ast}$, $\delta=\gamma^{\ast}$, $\nu=\mu^{\ast}$ and $\eta=\xi^{\ast}$, the system of equations (2.340), (2.341), (2.384) and (2.385) admits the complex conjugation reduction; that is, the reduction $r_{n}=\sigma q_{n}^{\ast}$, $v_{n}=\sigma u_{n}^{\ast}$ with a real constant $\sigma$ on the right-hand sides of the equations implies the reduction $\widetilde{r}_{n}=\sigma\widetilde{q}_{n}^{\,\ast}$, $v_{n+1}=\sigma u_{n+1}^{\,\ast}$ on the left-hand sides of the equations. However, this reduction is outside the scope of this paper.
If the parameters satisfy the conditions:
$\gamma=\alpha^{\ast},\hskip 14.22636pt\delta=\beta^{\ast},\hskip 14.22636pt\xi=\mu^{\ast},\hskip 14.22636pt\eta=\nu^{\ast},$
the fully discrete system of equations (2.340)–(2.343) admits the Hermitian conjugation reduction; that is, the reduction $r_{n}=q_{n}^{\dagger}$, $v_{n}=u_{n}^{\dagger}$ on the right-hand sides of the equations implies the reduction $\widetilde{r}_{n}=\widetilde{q}_{n}^{\,\dagger}$, $v_{n+1}=u_{n+1}^{\,\dagger}$ on the left-hand sides of the equations. In particular, if we set
$\alpha=-\gamma=2\mathrm{i}/h,\hskip 14.22636pt\beta=\delta=1,\hskip 14.22636pt\mu=\xi=1,\hskip 14.22636pt\nu=-\eta=2\mathrm{i}/\varDelta,\hskip 14.22636pth,\varDelta\in\mathbb{R}_{\neq 0},$ (2.386)
and impose this reduction, we obtain
$\widetilde{q}_{n}=$ $\;\left(I-\frac{\mathrm{i}h}{2}u_{n}u_{n}^{\dagger}\right)^{-1}\left[\left(1+\frac{h\varDelta}{4}\right)q_{n}+\mathrm{i}hu_{n}-\frac{\mathrm{i}h}{2}\left(1-\frac{h\varDelta}{4}\right)u_{n}u_{n}^{\dagger}q_{n}\right]$
×$\times\left[\left(1-\frac{h\varDelta}{4}\right)I+\frac{\mathrm{i}h}{2}\left(1+\frac{h\varDelta}{4}\right)u_{n}^{\dagger}u_{n}+\frac{h\varDelta}{2}u_{n}^{\dagger}q_{n}\right]^{-1}\left(I-\frac{\mathrm{i}h}{2}u_{n}^{\dagger}u_{n}\right),$ (2.387)
$u_{n+1}=$ $\;\left(I-\frac{\mathrm{i}\varDelta}{2}q_{n}q_{n}^{\dagger}\right)^{-1}\left[\left(1+\frac{h\varDelta}{4}\right)u_{n}-\mathrm{i}\varDelta q_{n}+\frac{\mathrm{i}\varDelta}{2}\left(1-\frac{h\varDelta}{4}\right)q_{n}q_{n}^{\dagger}u_{n}\right]$
×$\times\left[\left(1-\frac{h\varDelta}{4}\right)I-\frac{\mathrm{i}\varDelta}{2}\left(1+\frac{h\varDelta}{4}\right)q_{n}^{\dagger}q_{n}+\frac{h\varDelta}{2}q_{n}^{\dagger}u_{n}\right]^{-1}\left(I-\frac{\mathrm{i}\varDelta}{2}q_{n}^{\dagger}q_{n}\right),$ (2.388)
which can be simplified to
$\left\{\begin{split}&\widetilde{q}_{n}=\frac{\left(1+\frac{h\varDelta}{4}\right)q_{n}+\mathrm{i}hu_{n}-\frac{\mathrm{i}h}{2}\left(1-\frac{h\varDelta}{4}\right)|u_{n}|^{2}q_{n}}{\left(1-\frac{h\varDelta}{4}\right)+\frac{\mathrm{i}h}{2}\left(1+\frac{h\varDelta}{4}\right)|u_{n}|^{2}+\frac{h\varDelta}{2}u_{n}^{\ast}q_{n}},\\[4.2679pt] &u_{n+1}=\frac{\left(1+\frac{h\varDelta}{4}\right)u_{n}-\mathrm{i}\varDelta q_{n}+\frac{\mathrm{i}\varDelta}{2}\left(1-\frac{h\varDelta}{4}\right)|q_{n}|^{2}u_{n}}{\left(1-\frac{h\varDelta}{4}\right)-\frac{\mathrm{i}\varDelta}{2}\left(1+\frac{h\varDelta}{4}\right)|q_{n}|^{2}+\frac{h\varDelta}{2}q_{n}^{\ast}u_{n}},\end{split}\right.$ (2.389)
in the case of the scalar dependent variables. Note that (2.389) can be rewritten as
$\left\{\begin{split}&\mathrm{i}\frac{\widetilde{q}_{n}-q_{n}}{h}=\frac{\frac{\mathrm{i}\varDelta}{2}q_{n}-u_{n}+|u_{n}|^{2}q_{n}-\frac{\mathrm{i}\varDelta}{2}u_{n}^{\ast}q_{n}^{2}}{\left(1-\frac{h\varDelta}{4}\right)+\frac{\mathrm{i}h}{2}\left(1+\frac{h\varDelta}{4}\right)|u_{n}|^{2}+\frac{h\varDelta}{2}u_{n}^{\ast}q_{n}},\\[4.2679pt] &\mathrm{i}\frac{u_{n+1}-u_{n}}{\varDelta}=\frac{\frac{\mathrm{i}h}{2}u_{n}+q_{n}-|q_{n}|^{2}u_{n}-\frac{\mathrm{i}h}{2}q_{n}^{\ast}u_{n}^{2}}{\left(1-\frac{h\varDelta}{4}\right)-\frac{\mathrm{i}\varDelta}{2}\left(1+\frac{h\varDelta}{4}\right)|q_{n}|^{2}+\frac{h\varDelta}{2}q_{n}^{\ast}u_{n}},\end{split}\right.$ (2.390)
which can be interpreted as a fully discrete analog of the massive Thirring model in light-cone coordinates (2.16). We mention that another discrete analog of the massive Thirring model in light-cone coordinates was proposed in [35].
In the continuous limit of time ($h\to 0$), the fully discrete massive Thirring model (2.390) reduces to the semi-discrete massive Thirring model:
$\left\{\begin{split}&\mathrm{i}q_{n,t}=\frac{\mathrm{i}\varDelta}{2}q_{n}-u_{n}+|u_{n}|^{2}q_{n}-\frac{\mathrm{i}\varDelta}{2}u_{n}^{\ast}q_{n}^{2},\\[4.2679pt] &\mathrm{i}\frac{u_{n+1}-u_{n}}{\varDelta}=\frac{q_{n}-|q_{n}|^{2}u_{n}}{1-\frac{\mathrm{i}\varDelta}{2}|q_{n}|^{2}},\end{split}\right.$ (2.391)
which is an integrable space-discretization of the massive Thirring model in light-cone coordinates (2.16); with minor manipulation, (2.391) can be identified with the semi-discrete massive Thirring model proposed in our previous paper (see (3.3) in [30]).
3 Bäcklund–Darboux transformation
We present the binary Bäcklund–Darboux transformation for the discrete massive Thirring model given by equations (2.340)–(2.343); note that Xu and Pelinovsky [15] obtained some relevant results previously.
Because the Lax matrices (2.29) and (2.41) have their origin in the binary Bäcklund–Darboux transformation for the continuous Chen–Lee–Liu hierarchy, the binary Bäcklund–Darboux transformation for the discrete massive Thirring model naturally has the same form. We use the hat $\;\widehat{}\;$ to denote the Bäcklund–Darboux transformation:
$\left[\begin{array}[]{c}\widehat{\Psi}_{1,n}\\ \widehat{\Psi}_{2,n}\\ \end{array}\right]={\cal L}({\cal U}_{n},{\cal V}_{n};a,b,c,d)\left[\begin{array}[]{c}\Psi_{1,n}\\ \Psi_{2,n}\\ \end{array}\right],$ (3.1)
where the Darboux matrix ${\cal L}({\cal U}_{n},{\cal V}_{n};a,b,c,d)$ is defined exactly in the same manner as in (2.29) and (2.41):
${\cal L}({\cal U}_{n},{\cal V}_{n};a,b,c,d)$
$:=\left[\begin{array}[]{cc}aI&O\\ O&cI\\ \end{array}\right]+\frac{ab-cd}{c\zeta^{2}+b}\left[\begin{array}[]{cc}{\cal U}_{n}{\cal V}_{n}(cdI-{\cal U}_{n}{\cal V}_{n})^{-1}&-c\zeta{\cal U}_{n}(cdI-{\cal V}_{n}{\cal U}_{n})^{-1}\\ -c\zeta{\cal V}_{n}(cdI-{\cal U}_{n}{\cal V}_{n})^{-1}&c^{2}\zeta^{2}(cdI-{\cal V}_{n}{\cal U}_{n})^{-1}\\ \end{array}\right]$ (3.6)
$\hphantom{:}=\left[\begin{array}[]{cc}aI&O\\ O&cI\\ \end{array}\right]+\frac{ab-cd}{c\zeta^{2}+b}\left[\begin{array}[]{c}{\cal U}_{n}\\ -c\zeta I\\ \end{array}\right](cdI-{\cal V}_{n}{\cal U}_{n})^{-1}\left[\begin{array}[]{cc}{\cal V}_{n}&-c\zeta I\\ \end{array}\right].$ (3.12)
Here, $a$, $b$, $c$ and $d$ are nonzero constant parameters satisfying the condition $ab-cd\neq 0$. Similarly to (2.53), ${\cal L}({\cal U}_{n},{\cal V}_{n};a,b,c,d)$ can be rewritten up to an overall factor as
$\frac{\zeta+\frac{b}{c\zeta}}{\frac{a}{d}\zeta+\frac{1}{\zeta}}{\cal L}({\cal U}_{n},{\cal V}_{n};a,b,c,d)$
$\hphantom{:}=\left[\begin{array}[]{cc}dI&O\\ O&bI\\ \end{array}\right]+\frac{ab-cd}{a+\frac{d}{\zeta^{2}}}\left[\begin{array}[]{c}-\frac{d}{\zeta}I\\ {\cal V}_{n}\\ \end{array}\right](cdI-{\cal U}_{n}{\cal V}_{n})^{-1}\left[\begin{array}[]{cc}-\frac{d}{\zeta}I&{\cal U}_{n}\\ \end{array}\right].$ (3.18)
The definition (3.12) and the alternative expression (3.18) together with the linear problem (3.1) imply that the unknown functions ${\cal U}_{n}$ and ${\cal V}_{n}$ can be expressed in terms of the linear eigenfunction as
${\cal U}_{n}=-a\left.(\zeta\Psi_{1,n}^{(1)}\Psi_{2,n}^{(1)\,-1})\right|_{\zeta^{2}=-\frac{d}{a}},\hskip 14.22636pt{\cal V}_{n}=c\left.(\zeta\Psi_{2,n}^{(2)}\Psi_{1,n}^{(2)\,-1})\right|_{\zeta^{2}=-\frac{b}{c}},$ (3.19)
respectively. The superscript ${}^{(j)}\;(j=1,2)$ of the linear eigenfunction is used to stress that we can choose different linear eigenfunctions to define ${\cal U}_{n}$ and ${\cal V}_{n}$. Note that with an appropriate choice of the boundary/initial conditions, the functions $\zeta\Psi_{1,n}^{(1)}\Psi_{2,n}^{(1)\,-1}$ and $\zeta\Psi_{2,n}^{(2)}\Psi_{1,n}^{(2)\,-1}$ can be chosen as even functions in $\zeta$ (see (2.18) with (2.29) or (2.30) with (2.41)).
The compatibility condition for the overdetermined linear equations (2.18) and (3.1) is given by
${\cal L}(\widehat{q}_{n},\widehat{r}_{n};\mu,\nu,\xi,\eta){\cal L}({\cal U}_{n},{\cal V}_{n};a,b,c,d)={\cal L}({\cal U}_{n+1},{\cal V}_{n+1};a,b,c,d){\cal L}(q_{n},r_{n};\mu,\nu,\xi,\eta),$ (3.20)
while the compatibility condition for the overdetermined linear equations (2.30) and (3.1) is given by
${\cal L}(\widetilde{{\cal U}}_{n},\widetilde{{\cal V}}_{n};a,b,c,d){\cal L}(u_{n},v_{n};\alpha,\beta,\gamma,\delta)={\cal L}(\widehat{u}_{n},\widehat{v}_{n};\alpha,\beta,\gamma,\delta){\cal L}({\cal U}_{n},{\cal V}_{n};a,b,c,d).$ (3.21)
Here, we assume that $\nu/\xi$, $\eta/\mu$, $\beta/\gamma$, $\delta/\alpha$, $b/c$ and $d/a$ are pairwise distinct.
Because equations (3.20) and (3.21) have the same form as the discrete zero-curvature condition (2.42), we can solve these equations in the same way as in subsection 2.3. Thus, we can express $\widehat{q}_{n}$, $\widehat{r}_{n}$, $\widehat{u}_{n}$ and $\widehat{v}_{n}$ in terms of $q_{n}$, $r_{n}$, $u_{n}$, $v_{n}$, ${\cal U}_{n}$ and ${\cal V}_{n}$. The expressions are given by equations (2.340)–(2.343) with an appropriate replacement of the parameters, variables and notation, i.e.,
$\widehat{q}_{n}=$ $\;(cdI-{\cal U}_{n}{\cal V}_{n})^{-1}\left[-ad(b\mu-c\eta)q_{n}+\mu\eta(ab-cd){\cal U}_{n}-(a\eta-d\mu){\cal U}_{n}{\cal V}_{n}q_{n}\right]$
×$\times\left[bc(a\eta-d\mu)I+(b\mu-c\eta){\cal V}_{n}{\cal U}_{n}-(ab-cd){\cal V}_{n}q_{n}\right]^{-1}(cdI-{\cal V}_{n}{\cal U}_{n}),$ (3.22)
$\widehat{r}_{n}=$ $\;(abI-{\cal V}_{n}{\cal U}_{n})\left[ad(b\xi-c\nu)I+(a\nu-d\xi){\cal V}_{n}{\cal U}_{n}-(ab-cd)r_{n}{\cal U}_{n}\right]^{-1}$
×$\times\left[-bc(a\nu-d\xi)r_{n}+\nu\xi(ab-cd){\cal V}_{n}-(b\xi-c\nu)r_{n}{\cal U}_{n}{\cal V}_{n}\right](abI-{\cal U}_{n}{\cal V}_{n})^{-1},$ (3.23)
$\widehat{u}_{n}=$ $\;(cdI-{\cal U}_{n}{\cal V}_{n})^{-1}\left[ad(\alpha b-\delta c)u_{n}-\alpha\delta(ab-cd){\cal U}_{n}-(\alpha d-\delta a){\cal U}_{n}{\cal V}_{n}u_{n}\right]$
×$\times\left[bc(\alpha d-\delta a)I-(\alpha b-\delta c){\cal V}_{n}{\cal U}_{n}+(ab-cd){\cal V}_{n}u_{n}\right]^{-1}(cdI-{\cal V}_{n}{\cal U}_{n}),$ (3.24)
$\widehat{v}_{n}=$ $\;(abI-{\cal V}_{n}{\cal U}_{n})\left[ad(\beta c-\gamma b)I-(\beta a-\gamma d){\cal V}_{n}{\cal U}_{n}+(ab-cd)v_{n}{\cal U}_{n}\right]^{-1}$
×$\times\left[bc(\beta a-\gamma d)v_{n}-\beta\gamma(ab-cd){\cal V}_{n}-(\beta c-\gamma b)v_{n}{\cal U}_{n}{\cal V}_{n}\right](abI-{\cal U}_{n}{\cal V}_{n})^{-1}.$ (3.25)
As discussed in subsection 2.4, if the parameters satisfy the conditions:
$\gamma=\alpha^{\ast},\hskip 14.22636pt\delta=\beta^{\ast},\hskip 14.22636pt\xi=\mu^{\ast},\hskip 14.22636pt\eta=\nu^{\ast},$
the fully discrete system of equations (2.340)–(2.343) admits the Hermitian conjugation reduction; the Bäcklund–Darboux transformation formulas (3.22)–(3.25) are compatible with this reduction if we further impose the conditions:
$c=a^{\ast},\hskip 14.22636ptd=b^{\ast},\hskip 14.22636pt{\cal V}_{n}={\cal U}_{n}^{\dagger}.$
That is, the reduction $r_{n}=q_{n}^{\dagger}$, $v_{n}=u_{n}^{\dagger}$ on the right-hand sides of the equations implies the reduction $\widehat{r}_{n}=\widehat{q}_{n}^{\,\dagger}$, $\widehat{v}_{n}=\widehat{u}_{n}^{\,\dagger}$ on the left-hand sides of the equations.
In the remaining part of this section, we consider the case of the scalar dependent variables and choose the values of the parameters as given in (2.386). Thus, the Hermitian conjugation reduction is simplified to the complex conjugation reduction; the Lax-pair representation and the formulas for the Bäcklund–Darboux transformation can be summarized as follows.
The compatibility condition for the overdetermined linear equations:
$\left[\begin{array}[]{c}\Psi_{1,n+1}\\ \Psi_{2,n+1}\\ \end{array}\right]=\left\{I+\frac{\mathrm{i}\varDelta}{\left(1-\frac{\mathrm{i}\varDelta}{2}\zeta^{2}\right)\left(1-\frac{\mathrm{i}\varDelta}{2}|q_{n}|^{2}\right)}\left[\begin{array}[]{cc}|q_{n}|^{2}&-\zeta q_{n}\\ -\zeta q_{n}^{\ast}&\zeta^{2}\\ \end{array}\right]\right\}\left[\begin{array}[]{c}\Psi_{1,n}\\ \Psi_{2,n}\\ \end{array}\right],$ (3.26)
and
$\left[\begin{array}[]{c}\widetilde{\Psi}_{1,n}\\ \widetilde{\Psi}_{2,n}\\ \end{array}\right]=\left\{I+\frac{\mathrm{i}h}{\left(1-\frac{\mathrm{i}h}{2}\frac{1}{\zeta^{2}}\right)\left(1-\frac{\mathrm{i}h}{2}|u_{n}|^{2}\right)}\left[\begin{array}[]{cc}\frac{1}{\zeta^{2}}&-\frac{1}{\zeta}u_{n}\\[2.84526pt] -\frac{1}{\zeta}u_{n}^{\ast}&|u_{n}|^{2}\\ \end{array}\right]\right\}\left[\begin{array}[]{c}\Psi_{1,n}\\ \Psi_{2,n}\\ \end{array}\right],$ (3.27)
is equivalent to the fully discrete massive Thirring model in light-cone coordinates (2.389) (or (2.390)). Here, we rescaled the temporal Lax matrix as in (2.53). This Lax-pair representation is form-invariant under the binary Bäcklund–Darboux transformation:
$\left[\begin{array}[]{c}\widehat{\Psi}_{1,n}\\ \widehat{\Psi}_{2,n}\\ \end{array}\right]=\left\{\left[\begin{array}[]{cc}a&0\\ 0&a^{\ast}\\ \end{array}\right]+\frac{ab-a^{\ast}b^{\ast}}{(a^{\ast}\zeta^{2}+b)\left(a^{\ast}b^{\ast}-|{\cal U}_{n}|^{2}\right)}\left[\begin{array}[]{cc}|{\cal U}_{n}|^{2}&-a^{\ast}\zeta{\cal U}_{n}\\ -a^{\ast}\zeta{\cal U}_{n}^{\ast}&a^{*\hskip 1.0pt2}\zeta^{2}\\ \end{array}\right]\right\}\left[\begin{array}[]{c}\Psi_{1,n}\\ \Psi_{2,n}\\ \end{array}\right],$
where the unknown function ${\cal U}_{n}$ is expressed in terms of the linear eigenfunction as
${\cal U}_{n}=-a\left.(\zeta\Psi_{1,n}\Psi_{2,n}^{-1})\right|_{\zeta^{2}=-\frac{b^{\ast}}{a}},$ (3.28)
and the transformed variables are given by
$\widehat{q}_{n}=$ $\;\frac{-ab^{\ast}(2\mathrm{i}a^{\ast}\!+\varDelta b)q_{n}-2\mathrm{i}(ab-a^{\ast}b^{\ast}){\cal U}_{n}+(2\mathrm{i}a+\varDelta b^{\ast})|{\cal U}_{n}|^{2}q_{n}}{-a^{\ast}b(2\mathrm{i}a+\varDelta b^{\ast})+(2\mathrm{i}a^{\ast}\!+\varDelta b)|{\cal U}_{n}|^{2}-\varDelta(ab-a^{\ast}b^{\ast}){\cal U}_{n}^{\ast}q_{n}},$ (3.29)
$\widehat{u}_{n}=$ $\;\frac{ab^{\ast}(2\mathrm{i}b-ha^{\ast})u_{n}-2\mathrm{i}(ab-a^{\ast}b^{\ast}){\cal U}_{n}-(2\mathrm{i}b^{\ast}\!-ha)|{\cal U}_{n}|^{2}u_{n}}{a^{\ast}b(2\mathrm{i}b^{\ast}\!-ha)-(2\mathrm{i}b-ha^{\ast})|{\cal U}_{n}|^{2}+h(ab-a^{\ast}b^{\ast}){\cal U}_{n}^{\ast}u_{n}}.$ (3.30)
We start with the trivial zero solution $q_{n}=u_{n}=0$ for all time $m\in{\mathbb{Z}}$ and choose the solution of the linear equations (3.26) and (3.27) as
$\left[\begin{array}[]{c}\Psi_{1,n}(m)\\ \Psi_{2,n}(m)\\ \end{array}\right]=\left[\begin{array}[]{c}k\zeta\left(\frac{1+\frac{\mathrm{i}h}{2}\frac{1}{\zeta^{2}}}{1-\frac{\mathrm{i}h}{2}\frac{1}{\zeta^{2}}}\right)^{m}\\[8.53581pt] \left(\frac{1+\frac{\mathrm{i}\varDelta}{2}\zeta^{2}}{1-\frac{\mathrm{i}\varDelta}{2}\zeta^{2}}\right)^{n}\\ \end{array}\right],$
where $k$ is a nonzero complex constant. For convenience, we set $k=\kappa a^{\ast}/b^{\ast}$, where $\kappa$ is a nonzero complex constant. Then, (3.28) implies that
${\cal U}_{n}=\kappa a^{\ast}\left(\frac{1+\frac{\mathrm{i}\varDelta b^{\ast}}{2a}}{1-\frac{\mathrm{i}\varDelta b^{\ast}}{2a}}\right)^{n}\left(\frac{1-\frac{\mathrm{i}ha}{2b^{\ast}}}{1+\frac{\mathrm{i}ha}{2b^{\ast}}}\right)^{m},$
so we obtain from (3.29) and (3.30) the one-soliton solution of (2.389) (or (2.390)):
$\widehat{q}_{n}=$ $\;\frac{\kappa\left(1-\frac{a^{\ast}b^{\ast}}{ab}\right)\left(\frac{1+\frac{\mathrm{i}\varDelta b^{\ast}}{2a}}{1-\frac{\mathrm{i}\varDelta b^{\ast}}{2a}}\right)^{n}\left(\frac{1-\frac{\mathrm{i}ha}{2b^{\ast}}}{1+\frac{\mathrm{i}ha}{2b^{\ast}}}\right)^{m}}{\left(1-\frac{\mathrm{i}\varDelta b^{\ast}}{2a}\right)-\frac{a^{\ast}}{b}\left(1-\frac{\mathrm{i}\varDelta b}{2a^{\ast}}\right)|\kappa|^{2}\left|\frac{1+\frac{\mathrm{i}\varDelta b^{\ast}}{2a}}{1-\frac{\mathrm{i}\varDelta b^{\ast}}{2a}}\right|^{2n}\left|\frac{1-\frac{\mathrm{i}ha}{2b^{\ast}}}{1+\frac{\mathrm{i}ha}{2b^{\ast}}}\right|^{2m}},$
$\widehat{u}_{n}=$ $\;\frac{\kappa\left(1-\frac{a^{\ast}b^{\ast}}{ab}\right)\left(\frac{1+\frac{\mathrm{i}\varDelta b^{\ast}}{2a}}{1-\frac{\mathrm{i}\varDelta b^{\ast}}{2a}}\right)^{n}\left(\frac{1-\frac{\mathrm{i}ha}{2b^{\ast}}}{1+\frac{\mathrm{i}ha}{2b^{\ast}}}\right)^{m}}{-\frac{b^{\ast}}{a}\left(1+\frac{\mathrm{i}ha}{2b^{\ast}}\right)+\left(1+\frac{\mathrm{i}ha^{\ast}}{2b}\right)|\kappa|^{2}\left|\frac{1+\frac{\mathrm{i}\varDelta b^{\ast}}{2a}}{1-\frac{\mathrm{i}\varDelta b^{\ast}}{2a}}\right|^{2n}\left|\frac{1-\frac{\mathrm{i}ha}{2b^{\ast}}}{1+\frac{\mathrm{i}ha}{2b^{\ast}}}\right|^{2m}}.$
4 Yang–Baxter map
The discrete Lax matrix ${\cal L}(q_{n},r_{n};\mu,\nu,\xi,\eta)$ defined in (2.29) has the property:
${\cal L}(q_{n},r_{n};\mu,\nu,\xi,\eta){\cal L}(q_{n},r_{n};\xi,\eta,\mu,\nu)=\mu\xi I.$ (4.1)
Similarly, the discrete Lax matrix ${\cal L}(u_{n},v_{n};\alpha,\beta,\gamma,\delta)$ defined in (2.41) has the property:
${\cal L}(u_{n},v_{n};\alpha,\beta,\gamma,\delta){\cal L}(u_{n},v_{n};\gamma,\delta,\alpha,\beta)=\alpha\gamma I.$ (4.2)
Using the property (4.1), we can rewrite the discrete zero-curvature condition (2.42) as
${\cal L}(\widetilde{q}_{n},\widetilde{r}_{n};\xi,\eta,\mu,\nu){\cal L}(u_{n+1},v_{n+1};\alpha,\beta,\gamma,\delta)={\cal L}(u_{n},v_{n};\alpha,\beta,\gamma,\delta){\cal L}(q_{n},r_{n};\xi,\eta,\mu,\nu).$
Thus, the map $R_{\{\mu,\nu,\xi,\eta\}}^{\{\alpha,\beta,\gamma,\delta\}}:\ \left(q_{n},r_{n},u_{n+1},v_{n+1}\right)\mapsto\left(\widetilde{q}_{n},\widetilde{r}_{n},u_{n},v_{n}\right)$ takes exactly the same form as the map $\left(q_{n},r_{n},u_{n},v_{n}\right)\mapsto\left(\widetilde{q}_{n},\widetilde{r}_{n},u_{n+1},v_{n+1}\right)$ described by (2.340)–(2.343), but with the interchange of the parameters: $\mu\leftrightarrow\xi$ and $\nu\leftrightarrow\eta$. Hence, the map $R_{\{\mu,\nu,\xi,\eta\}}^{\{\alpha,\beta,\gamma,\delta\}}$ can be written explicitly as
$\widetilde{q}_{n}=$ $\;(\gamma\delta I-u_{n+1}v_{n+1})^{-1}\left[-\alpha\delta(\beta\xi-\gamma\nu)q_{n}+\nu\xi(\alpha\beta-\gamma\delta)u_{n+1}-(\alpha\nu-\delta\xi)u_{n+1}v_{n+1}q_{n}\right]$
×$\times\left[\beta\gamma(\alpha\nu-\delta\xi)I+(\beta\xi-\gamma\nu)v_{n+1}u_{n+1}-(\alpha\beta-\gamma\delta)v_{n+1}q_{n}\right]^{-1}(\gamma\delta I-v_{n+1}u_{n+1}),$ (4.3)
$\widetilde{r}_{n}=$ $\;(\alpha\beta I-v_{n+1}u_{n+1})\left[\alpha\delta(\beta\mu-\gamma\eta)I+(\alpha\eta-\delta\mu)v_{n+1}u_{n+1}-(\alpha\beta-\gamma\delta)r_{n}u_{n+1}\right]^{-1}$
×$\times\left[-\beta\gamma(\alpha\eta-\delta\mu)r_{n}+\mu\eta(\alpha\beta-\gamma\delta)v_{n+1}-(\beta\mu-\gamma\eta)r_{n}u_{n+1}v_{n+1}\right](\alpha\beta I-u_{n+1}v_{n+1})^{-1},$ (4.4)
$u_{n}=$ $\;(\mu\nu I-q_{n}r_{n})^{-1}\left[\nu\xi(\alpha\eta-\delta\mu)u_{n+1}+\alpha\delta(\mu\nu-\xi\eta)q_{n}-(\alpha\nu-\delta\xi)q_{n}r_{n}u_{n+1}\right]$
×$\times\left[\mu\eta(\alpha\nu-\delta\xi)I-(\alpha\eta-\delta\mu)r_{n}q_{n}-(\mu\nu-\xi\eta)r_{n}u_{n+1}\right]^{-1}(\mu\nu I-r_{n}q_{n}),$ (4.5)
$v_{n}=$ $\;(\xi\eta I-r_{n}q_{n})\left[\nu\xi(\beta\mu-\gamma\eta)I-(\beta\xi-\gamma\nu)r_{n}q_{n}-(\mu\nu-\xi\eta)v_{n+1}q_{n}\right]^{-1}$
×$\times\left[\mu\eta(\beta\xi-\gamma\nu)v_{n+1}+\beta\gamma(\mu\nu-\xi\eta)r_{n}-(\beta\mu-\gamma\eta)v_{n+1}q_{n}r_{n}\right](\xi\eta I-q_{n}r_{n})^{-1}.$ (4.6)
As stated in subsection 2.2, the nonzero constant parameters $\mu$, $\nu$, $\xi$, $\eta$, $\alpha$, $\beta$, $\gamma$ and $\delta$ are required to satisfy the conditions $\mu\nu-\xi\eta\neq 0$, $\alpha\beta-\gamma\delta\neq 0$, $\alpha\nu-\delta\xi\neq 0$, $\alpha\eta-\delta\mu\neq 0$, $\beta\mu-\gamma\eta\neq 0$ and $\beta\xi-\gamma\nu\neq 0$. In other words, the values of $\nu/\xi$, $\eta/\mu$, $\beta/\gamma$ and $\delta/\alpha$ are pairwise distinct.
$(q_{n},r_{n};\mu,\nu,\xi,\eta)$$(\widetilde{q}_{n},\widetilde{r}_{n};\mu,\nu,\xi,\eta)$$(u_{n},v_{n};\alpha,\beta,\gamma,\delta)$$(u_{n+1},v_{n+1};\alpha,\beta,\gamma,\delta)$$R_{\{\mu,\nu,\xi,\eta\}}^{\{\alpha,\beta,\gamma,\delta\}}$
The map $R_{\{\mu,\nu,\xi,\eta\}}^{\{\alpha,\beta,\gamma,\delta\}}$ described by (4.3)–(4.6) is a parameter-dependent Yang–Baxter map because its admits the Lax representation [36, 10] given by the discrete zero-curvature condition (2.42). If we adopt the original definition by Veselov and coworkers [39, 37, 3, 38], the Lax representation for this Yang–Baxter map is given by taking the matrix transpose (or matrix inverse) of (2.42). For the Lax representation to be proper, we have to confirm the uniqueness of factorization [2, 39] for the product of three Lax matrices [36, 10]. In our case, assume that the relation
${\cal L}(q^{\prime},r^{\prime};\mu,\nu,\xi,\eta){\cal L}(u^{\prime},v^{\prime};\alpha,\beta,\gamma,\delta){\cal L}({\cal U}^{\prime},{\cal V}^{\prime};a,b,c,d)$
$={\cal L}(q,r;\mu,\nu,\xi,\eta){\cal L}(u,v;\alpha,\beta,\gamma,\delta){\cal L}({\cal U},{\cal V};a,b,c,d),$ (4.7)
holds, where the Lax matrices are given by (2.29), (2.41) and (3.12), all the parameters are nonzero and $\nu/\xi$, $\eta/\mu$, $\beta/\gamma$, $\delta/\alpha$, $b/c$ and $d/a$ are pairwise distinct. Then, we should have $q^{\prime}=q$, $r^{\prime}=r$, $u^{\prime}=u$, $v^{\prime}=v$, ${\cal U}^{\prime}={\cal U}$ and ${\cal V}^{\prime}={\cal V}$. That is, the equality (4.7) holds only in the trivial case.
This uniqueness of factorization can be checked as follows. With the aid of the relations (4.1), (4.2) and
${\cal L}({\cal U},{\cal V};a,b,c,d){\cal L}({\cal U},{\cal V};c,d,a,b)=acI,$
we can rewrite (4.7) as
${\cal L}(q,r;\xi,\eta,\mu,\nu){\cal L}(q^{\prime},r^{\prime};\mu,\nu,\xi,\eta)$
$=\frac{\mu\xi}{ac\alpha\gamma}{\cal L}(u,v;\alpha,\beta,\gamma,\delta){\cal L}({\cal U},{\cal V};a,b,c,d){\cal L}({\cal U}^{\prime},{\cal V}^{\prime};c,d,a,b){\cal L}(u^{\prime},v^{\prime};\gamma,\delta,\alpha,\beta).$ (4.8)
Using (2.29), we can expand the left-hand side of (4.8) as
$\left[\begin{array}[]{cc}\mu\xi I&O\\ O&\mu\xi I\\ \end{array}\right]-\frac{\mu\nu-\xi\eta}{\mu\zeta^{2}+\eta}\left[\begin{array}[]{c}q\\ -\mu\zeta I\\ \end{array}\right](\mu\nu I-rq)^{-1}\left[\begin{array}[]{cc}\mu r&-\mu\xi\zeta I\\ \end{array}\right]$ (4.14)
$+\frac{\mu\nu-\xi\eta}{\xi\zeta^{2}+\nu}\left[\begin{array}[]{c}\xi q^{\prime}\\ -\mu\xi\zeta I\\ \end{array}\right](\xi\eta I-r^{\prime}q^{\prime})^{-1}\left[\begin{array}[]{cc}r^{\prime}&-\xi\zeta I\\ \end{array}\right]$ (4.18)
$-(\mu\nu-\xi\eta)\left(\frac{\mu}{\mu\zeta^{2}+\eta}-\frac{\xi}{\xi\zeta^{2}+\nu}\right)\left[\begin{array}[]{c}q\\ -\mu\zeta I\\ \end{array}\right](\mu\nu I-rq)^{-1}rq^{\prime}(\xi\eta I-r^{\prime}q^{\prime})^{-1}\left[\begin{array}[]{cc}r^{\prime}&-\xi\zeta I\\ \end{array}\right]$ (4.22)
$+(\mu\nu-\xi\eta)\left(\frac{\eta}{\mu\zeta^{2}+\eta}-\frac{\nu}{\xi\zeta^{2}+\nu}\right)\mu\xi\left[\begin{array}[]{c}q\\ -\mu\zeta I\\ \end{array}\right](\mu\nu I-rq)^{-1}(\xi\eta I-r^{\prime}q^{\prime})^{-1}\left[\begin{array}[]{cc}r^{\prime}&-\xi\zeta I\\ \end{array}\right].$ (4.26)
By multiplying (4.8) both from the left and right by the diagonal matrix:
$\left[\begin{array}[]{cc}I&O\\ O&\frac{1}{\zeta}I\\ \end{array}\right],$
and taking the residue at $\zeta^{2}=-\eta/\mu$ and $\zeta^{2}=-\nu/\xi$, respectively, we have $r^{\prime}=r$ and $q^{\prime}=q$. Repeating a similar procedure, we also have $v^{\prime}=v$, $u^{\prime}=u$, ${\cal V}^{\prime}={\cal V}$ and ${\cal U}^{\prime}={\cal U}$.
If the parameters satisfy the conditions discussed in subsection 2.4:
$\gamma=\alpha^{\ast},\hskip 14.22636pt\delta=\beta^{\ast},\hskip 14.22636pt\xi=\mu^{\ast},\hskip 14.22636pt\eta=\nu^{\ast},$
the Yang–Baxter map described by (4.3)–(4.6) admits the Hermitian conjugation reduction; that is, the reduction $r_{n}=q_{n}^{\dagger}$, $v_{n+1}=u_{n+1}^{\dagger}$ on the right-hand sides of the equations implies the reduction $\widetilde{r}_{n}=\widetilde{q}_{n}^{\,\dagger}$, $v_{n}=u_{n}^{\,\dagger}$ on the left-hand sides of the equations. Thus, we obtain the reduced form of the Yang–Baxter map $R_{\{\mu,\nu\}}^{\{\alpha,\beta\}}:\ \left(q_{n},u_{n+1}\right)\mapsto\left(\widetilde{q}_{n},u_{n}\right)$ as given by
$\widetilde{q}_{n}=$ $\left(\alpha^{\ast}\beta^{\ast}I-u_{n+1}u_{n+1}^{\dagger}\right)^{-1}\left[-\alpha\beta^{\ast}(\beta\mu^{\ast}-\alpha^{\ast}\nu)q_{n}+\nu\mu^{\ast}(\alpha\beta-\alpha^{\ast}\beta^{\ast})u_{n+1}\right.$
$\left.\mbox{}-(\alpha\nu-\beta^{\ast}\mu^{\ast})u_{n+1}u_{n+1}^{\dagger}q_{n}\right]\left[\beta\alpha^{\ast}(\alpha\nu-\beta^{\ast}\mu^{\ast})I+(\beta\mu^{\ast}-\alpha^{\ast}\nu)u_{n+1}^{\dagger}u_{n+1}\right.$
$\left.\mbox{}-(\alpha\beta-\alpha^{\ast}\beta^{\ast})u_{n+1}^{\dagger}q_{n}\right]^{-1}\left(\alpha^{\ast}\beta^{\ast}I-u_{n+1}^{\dagger}u_{n+1}\right),$ (4.27)
$u_{n}=$ $\left(\mu\nu I-q_{n}q_{n}^{\dagger}\right)^{-1}\left[\nu\mu^{\ast}(\alpha\nu^{\ast}-\beta^{\ast}\mu)u_{n+1}+\alpha\beta^{\ast}(\mu\nu-\mu^{\ast}\nu^{\ast})q_{n}-(\alpha\nu-\beta^{\ast}\mu^{\ast})q_{n}q_{n}^{\dagger}u_{n+1}\right]$
×$\times\left[\mu\nu^{\ast}(\alpha\nu-\beta^{\ast}\mu^{\ast})I-(\alpha\nu^{\ast}-\beta^{\ast}\mu)q_{n}^{\dagger}q_{n}-(\mu\nu-\mu^{\ast}\nu^{\ast})q_{n}^{\dagger}u_{n+1}\right]^{-1}\left(\mu\nu I-q_{n}^{\dagger}q_{n}\right).$ (4.28)
$(q_{n};\mu,\nu)$$(\widetilde{q}_{n};\mu,\nu)$$(u_{n};\alpha,\beta)$$(u_{n+1};\alpha,\beta)$$R_{\{\mu,\nu\}}^{\{\alpha,\beta\}}$
If we set the values of the parameters as in (2.386), i.e.,
$\alpha=-\gamma=2\mathrm{i}/h,\hskip 14.22636pt\beta=\delta=1,\hskip 14.22636pt\mu=\xi=1,\hskip 14.22636pt\nu=-\eta=2\mathrm{i}/\varDelta,\hskip 14.22636pth,\varDelta\in\mathbb{R}_{\neq 0},$
and consider the simplest case of the scalar dependent variables, we obtain the map $\left(q_{n},u_{n+1}\right)\mapsto\left(\widetilde{q}_{n},u_{n}\right)$ in the form:
$\left\{\begin{split}&\widetilde{q}_{n}=\frac{\left(1-\frac{h\varDelta}{4}\right)q_{n}+\mathrm{i}hu_{n+1}-\frac{\mathrm{i}h}{2}\left(1+\frac{h\varDelta}{4}\right)|u_{n+1}|^{2}q_{n}}{\left(1+\frac{h\varDelta}{4}\right)+\frac{\mathrm{i}h}{2}\left(1-\frac{h\varDelta}{4}\right)|u_{n+1}|^{2}-\frac{h\varDelta}{2}u_{n+1}^{\ast}q_{n}},\\[4.2679pt] &u_{n}=\frac{\left(1-\frac{h\varDelta}{4}\right)u_{n+1}+\mathrm{i}\varDelta q_{n}-\frac{\mathrm{i}\varDelta}{2}\left(1+\frac{h\varDelta}{4}\right)|q_{n}|^{2}u_{n+1}}{\left(1+\frac{h\varDelta}{4}\right)+\frac{\mathrm{i}\varDelta}{2}\left(1-\frac{h\varDelta}{4}\right)|q_{n}|^{2}-\frac{h\varDelta}{2}q_{n}^{\ast}u_{n+1}}.\end{split}\right.$ (4.29)
Note that (4.29) can be rewritten as
$\left\{\begin{split}&\mathrm{i}\frac{\widetilde{q}_{n}-q_{n}}{h}=\frac{-\frac{\mathrm{i}\varDelta}{2}q_{n}-u_{n+1}+|u_{n+1}|^{2}q_{n}+\frac{\mathrm{i}\varDelta}{2}u_{n+1}^{\ast}q_{n}^{2}}{\left(1+\frac{h\varDelta}{4}\right)+\frac{\mathrm{i}h}{2}\left(1-\frac{h\varDelta}{4}\right)|u_{n+1}|^{2}-\frac{h\varDelta}{2}u_{n+1}^{\ast}q_{n}},\\[4.2679pt] &\mathrm{i}\frac{u_{n}-u_{n+1}}{\varDelta}=\frac{-\frac{\mathrm{i}h}{2}u_{n+1}-q_{n}+|q_{n}|^{2}u_{n+1}+\frac{\mathrm{i}h}{2}q_{n}^{\ast}u_{n+1}^{2}}{\left(1+\frac{h\varDelta}{4}\right)+\frac{\mathrm{i}\varDelta}{2}\left(1-\frac{h\varDelta}{4}\right)|q_{n}|^{2}-\frac{h\varDelta}{2}q_{n}^{\ast}u_{n+1}}.\end{split}\right.$ (4.30)
Clearly, (4.30) can be interpreted as a fully discrete analog of the massive Thirring model in light-cone coordinates (2.16), so the Yang–Baxter map $\left(q_{n},u_{n+1}\right)\mapsto\left(\widetilde{q}_{n},u_{n}\right)$ given by (4.27) and (4.28) indeed belongs to the class discussed in the introduction.
5 Concluding remarks
In this paper, we use a discrete Lax matrix representing the binary Bäcklund–Darboux transformation for the continuous derivative nonlinear Schrödinger (Chen–Lee–Liu[19]) hierarchy as the main building block. By solving the corresponding discrete zero-curvature condition, we obtain a fully discrete system, which provides, with a particular choice of the parameters, an integrable discretization of the massive Thirring model in light-cone (or characteristic) coordinates. The binary Bäcklund–Darboux transformation for the discrete massive Thirring model can be constructed in a straightforward manner and the one-soliton solution is presented explicitly. Using some symmetry properties of the discrete Lax matrix, we can rewrite the fully discrete system, which includes the discrete massive Thirring model as a particular case, as a parameter-dependent Yang–Baxter map. The Yang–Baxter map satisfies the parameter-dependent version [36, 10] of the quantum Yang–Baxter equation (1.1), i.e.,
$R_{12}(\mbox{\boldmath$\alpha$},\mbox{\boldmath$\beta$})R_{13}(\mbox{\boldmath$\alpha$},\mbox{\boldmath$\gamma$})R_{23}(\mbox{\boldmath$\beta$},\mbox{\boldmath$\gamma$})=R_{23}(\mbox{\boldmath$\beta$},\mbox{\boldmath$\gamma$})R_{13}(\mbox{\boldmath$\alpha$},\mbox{\boldmath$\gamma$})R_{12}(\mbox{\boldmath$\alpha$},\mbox{\boldmath$\beta$}).$
Thus, we obtain a new parameter-dependent Yang–Baxter map that admits a natural continuous limit for $R_{12}$; the two parameters therein can be understood as the lattice parameters. Note that this Yang–Baxter map can also be constructed using the definition of the unknown functions appearing in the discrete Lax matrix in terms of the linear eigenfunction (see (3.1) and (3.12) with (3.19)). It would be interesting to discuss how this Yang–Baxter map can be characterized within the framework of the matrix modified KP equation (see [11, 12] for the simpler case of the matrix KP equation).